• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Meekins

    Log entry time 08:24:16 on October 25,1999

    Entry number 29124

    keyword=Shift Summary

    Finnished taking data on P_r=400MeV.
    Took harp scan, carbon and 4cm dummy runs here for alignment. I think, that the harp scan was performed properly.

    We switched to the Pr=150 kinematics at 0300 and were down for 1.5 hr for a Hall B test of some sort.

    At 0515 TS0 died and would not respond to repeated attempts to reboot it from upstairs. We made an access and power cycled the crate 3 times. On the 3rd time the crate came back. We left the hall at 0600.

    Beam was restored at 0630
    Completed the production running on the P_r=150.

    Ended with target IOC problems